# Efficient Parallel Algorithms for Inpainting-Based Representations of 4K Images

### v1

@@ 4.1 CGNR Solver for Tonal Optimization
To apply $\bm{B}^{\top}\bm{B}$ to a vector $\bm{x}$ we first solve $\bm{A}\bm{y}=\bm{C}\bm{x}$ and then in the second step we solve $\bm{A}^{\top}\bm{z}=\bm{C}\bm{y}$. This leads to a nested problem structure, where the outer iterations optimize the color values at mask points, while the inner iterations solve inpainting problems with the matrix $\bm{A}$.
The linear system corresponding to the product $\bm{B}^{\top}\bm{w}$ involves the non-symmetric matrix $\bm{A}^{\top}$, which does not reduce to a symmetric system in the same way as $\bm{A}\bm{x}=\bm{b}$. Thus, we cannot solve it with the CG algorithm or our full multigrid ORAS method. As a remedy we use a symmetrized inpainting matrix $\tilde{\bm{A}}$ together with a modified $\tilde{\bm{C}}$, such that $\tilde{\bm{A}}\bm{u}=\tilde{\bm{C}}\bm{f}$ has the same solution as $\bm{A}\bm{u}=\bm{C}\bm{f}$. We can derive them with the following transformations, by using the fact that $\bm{C}\bm{u}=\bm{C}\bm{f}$, as the inpainting solution and the original image coincide at the mask pixels:
@@
We improve the nested CG approach in two ways. First, we replace the inner CG solvers for the two inpainting problems by our full multigrid ORAS method, which improves the runtime by more than a factor of 4. Secondly, we propose to replace the outer CG solver with CGNR, a variant of CG for the normal equations [10, 26, 60]. Besides better numerical stability compared to CG, CGNR also has the advantage that the original inpainting residual $\bm{r}=\bm{f}-\bm{A}\bm{u}$, and hence the MSE, is available in each iteration. Since we want to optimize the MSE, a stopping criterion based on the relative improvement of the MSE in an iteration is better suited for our purposes than the standard stopping criterion based on the $2$-norm of the residual of the normal equations $\|\bm{B}^{\top}(\bm{f}-\bm{B}\bm{g})\|_{2}$.
@@ 4.2 RAS Solver for Tonal Optimization
Taking into account the above, we solve the normal equations with a domain decomposition method instead of CGNR. We adapt a restricted additive Schwarz (RAS) method [13], which is similar to the one used in our fast inpainting solver. For more details on our domain decomposition method we refer to the first part of our companion paper [39]. Similar to the inpainting case, we subdivide the image domain $\Omega$ into $k\in\mathbb{N}$ overlapping subdomains $\Omega_{1},...,\Omega_{k}\subset\Omega$, such that $\cup_{i=1}^{k}\Omega_{i}=\Omega$. The resulting RAS tonal optimization is given in Algorithm 2. It consists of three steps that are iterated until the method is sufficiently converged.
@@
    $\bm{g}^{n+1}=\bm{g}^{n}+\sum_{i=1}^{k}\bm{R}_{i}^{\top}\bm{D}_{i}\bm{v}_{i}^{n}$ (10)
First we compute the global residual $\bm{r}^{n}$. This involves the application of the system matrix. As in Subsection 4.1 we evaluate the matrix-vector products without explicitly forming the matrix, by solving two inner inpainting problems. As the inpainting solution changes only slightly from one iteration to the next, it can be used as a good initialization for the next inpainting. This allows us to skip the initialization phase in our full multigrid ORAS inpainting and start directly with a V-cycle at the finest resolution level, which improves the runtime of both inpaintings significantly. This is an advantage of RAS compared to CGNR, where we apply the system matrix only to conjugate vectors, which change more from one iteration to the next, and thus cannot be used as a good initialization.
@@
Overall our RAS method consists of four types of problems, which are shown in Fig. 4. Our global tonal optimization problem is solved with RAS which includes two global inner inpainting problems per iteration to compute the global residual. They are solved with our multigrid ORAS inpainting method, proposed in the first part of our companion paper [39]. For each domain decomposition block in the RAS solver we also get local tonal optimization problems which we solve with CGNR. These again include two local inner inpainting problems per iteration that we also solve with our multigrid ORAS inpainting method.
For the domain decomposition we use blocks of size $64\times 64$, an overlap of $6$ pixels and a simple averaging for mask pixels in the overlap. In practice, this ensures sufficient communication between the blocks such that the method converges towards the global solution.
@@ 4.3 Voronoi Initialization
$\bm{g}_{i}=\bm{u}_{i}+\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}(\bm{f}_{j}-\bm{u}_{j})\quad\text{for all }i\in K,$ (11)
@@
$\bm{C}\bm{W}\bm{A}^{-1}\bm{C}\bm{g}=:\bm{C}\bm{W}\bm{u}=\bm{C}\bm{W}\bm{f}.$ (12)
We can e.g. solve it with a modified Richardson iteration:
$\bm{C}\bm{g}^{k+1}=\bm{C}\bm{g}^{k}+\tau(\bm{C}\bm{f}-\bm{C}\bm{W}\bm{A}^{-1}\bm{C}\bm{g}^{k}).$ (13)
@@ 5.1 Experimental Setup
For the experimental evaluation of our Delaunay densification spatial optimization method, we compare against the analytic approach (AA) by Belhachmi et al. [8], probabilistic sparsification (PS) [44], with (PS+NLPE) and without nonlocal pixel exchange, and the most recent neural approach by Schrader et al. [62]. The analytic approach is implemented with Floyd-Steinberg dithering [22] of the Laplacian magnitude. PS and NLPE are both implemented on the GPU and use our fast multigrid ORAS inpainting [39] to ensure fair comparisons. For PS we use candidate fractions $p=0.3$ and $q=0.005$, and NLPE is run for five cycles. As the neural approach requires more than 20 GB of graphics memory for runtime tests, we exclude them in our runtime comparison.
@@ 5.2 Spatial Optimization Experiments
In Fig. 7 we evaluate the impact of the choice of the initial mask on our Delaunay densification. We compare a uniform random initial mask with a Laplacian magnitude sampled initial mask. We use the same number of initial mask pixels and add a constant amount of new pixels in each iteration for both masks. We observe that for a small amount of iterations the Laplacian magnitude approach results in a significantly higher PSNR, while the difference becomes smaller the more iterations are used. This can be explained with the smaller number of initial mask pixels for larger numbers of iterations, which reduces the impact of the initial mask. It also shows that the improvement per iteration becomes quite small for iterations beyond $20$, while the additional runtime for each iteration is approximately constant as each requires exactly one inpainting. Thus $20$ iterations seem to be a good trade-off between runtime and reconstruction quality.
@@
Figure 7 shows the effect of increasing or decreasing the number of mask pixels that are added per iteration during the densification with a constant factor $t$. For a uniform random initial mask we observe that adding a constant amount of pixels in each iteration is not the optimal strategy. By increasing the number of added pixels by a factor of $t=1.08$ or $8$% in each iteration, we can improve the PSNR by $0.05$ dB. With a sampled Laplacian magnitude as an initial mask, however, the optimal factor is only $t=1.03$ and the improvement compared to a constant number of pixels is less than $0.01$ dB. These results also reinforce the conclusion that the Laplacian magnitude yields a suitable initial mask and suggests only a slightly increasing number of pixels added per iteration. Since the difference between adding a constant number of pixels and the optimal strategy is negligible, for simplicity we use a constant number of pixels for the rest of our experiments.
[…]
@@
We evaluated the runtimes in Fig. 7(b). While the inferior analytic approach is clearly the fastest, our method is the second fastest with an average runtime of around $0.4$ seconds. We achieve this by restricting our densification to $20$ iterations, which still produces masks of very high quality. Compared to PS and PS+NLPE we are several orders of magnitude faster, since they both require thousands of inpaintings. For the neural approach we cannot generate any runtime results on our GPU due to memory restrictions, but the runtimes reported by Schrader et al. [62] are quite similar to the results of our approach, albeit a bit faster for densities below $3$%. As they are obtained with a significantly more powerful GPU, our method would likely be faster on a similar GPU. Furthermore, our method has the advantage of needing significantly less GPU memory.
@@
The qualitative evaluation of the spatial optimization methods in Fig. 7(a) shows that our Delaunay densification approach (DD) consistently outperforms all its competitors on all tested mask densities, even with just $20$ iterations. It even surpasses the recent neural approach by Schrader et al. [62] by at least $0.75$ dB. Compared to the model-based competitors our method is especially well-suited for lower mask densities, which are the most interesting for inpainting-based compression. Here our method outperforms the analytic approach by more than $6$ dB, PS by more than $3.8$, and PS+NLPE by more than $2.6$ dB. This suggests that densification together with an adaptation to the local mask density with for example the Delaunay triangulation should be preferred over the probabilistic sparsification approach, as it significantly outperforms it in terms of quality as well as runtime.
@@ 5.3 Tonal Optimization Experiments
In order to quantify the effect of our Voronoi approach on the tonal optimization problem, we perform an experiment using our RAS tonal optimization method with and without the Voronoi initialization. Additionally, we compare against a CGNR solver with our ORAS multigrid inpainting. Fig. 11(b) shows the result for a $5$% mask density. For a better visualization, we have omitted the results for the CGNR solver with CG inpainting, as each iteration needs approximately $0.8$ seconds, which makes it clearly inferior to the other methods. We can observe that RAS converges significantly faster to the optimal PSNR compared to the CGNR method. When we use RAS together with the Voronoi initialization, we can stop the tonal optimization much earlier, even though the initial iteration takes longer. As the tonal optimization is a convex least squares problem, all tested methods eventually converge to the same optimal PSNR. Therefore, we omit a comparison of the reconstruction quality, as all methods end up with the same solution. The experiment also shows that the Voronoi initialization alone almost reaches the optimal PSNR up to $0.2$ dB. This proves that the Voronoi initialization is an essential tool for tonal optimization.
@@ 6 Conclusions and Outlook
Besides the computation of the inpainting itself, which we already tackled in the first part of our companion paper, inpainting-based compression also requires the optimization of the inpainting data. The optimization of the spatial data as well as the tonal data for homogeneous diffusion inpainting is a challenging task, and previous approaches were either slow or of inferior quality. As a remedy, we developed new methods for the data optimization of homogeneous diffusion inpainting that outperform previous approaches by a wide margin. Furthermore, most prior approaches could handle only relatively small images, while we are able to work with images of 4K resolution and higher with optimal linear resolution to runtime scaling. We achieved this by carefully adapting some of the most successful numerical concepts and transferring them to an image compression application. For our spatial optimization method we combined a mask densification approach with error-map dithering ideas using a Delaunay triangulation. For the tonal optimization we adopted a domain decomposition method that solves the corresponding normal equations in a matrix-free fashion. Additionally we proposed a Voronoi-based initialization strategy that allows us to stop the tonal optimization much earlier. This allows us to create high quality inpainting masks and optimized tonal values in a runtime of less than a second on a contemporary GPU. This constitutes a significant improvement over any comparable previous approaches and suggests a similar order of efficiency as JPEG2000, while solving a numerically much more challenging problem. Together with our real-time capable inpainting from the first part of our companion paper, we presented a basis for inpainting-based compression methods that may serve as serious alternatives to classical transform-based codecs, both in terms of quality as well as runtimes.
In our ongoing research we are working on extending our spatial and tonal optimization approaches to more sophisticated inpainting operators, such as anisotropic nonlinear diffusion. The latter has been shown to qualitatively outperform homogeneous diffusion.

### v2

@@ 5.1 CGNR Solver for Tonal Optimization
To compute the matrix-vector product $\bm{z}=\bm{B}^{\top}\bm{B}\bm{x}$ we first have to solve $\bm{A}\bm{y}=\bm{C}\bm{x}$ (for the step $\bm{y}=\bm{B}\bm{x}=\bm{A}^{-1}\bm{C}\bm{x}$), and then we have to solve $\bm{A}^{\top}\bm{C}\bm{z}=\bm{y}$ (for the step $\bm{z}=\bm{B}^{\top}\bm{y}=\bm{C}\bm{A}^{-\top}\bm{y}$). This leads to a nested problem structure with the outer tonal optimization problem and the two inner inpainting problems in each iteration.
The linear system corresponding to the product $\bm{B}^{\top}\bm{w}$ involves the non-symmetric matrix $\bm{A}^{\top}$, which does not reduce to a symmetric system in the same way as $\bm{A}\bm{x}=\bm{b}$. Thus, we cannot solve it with the CG algorithm or the multigrid ORAS inpainting solver. As a remedy we use a symmetrized inpainting matrix $\tilde{\bm{A}}$ together with a modified $\tilde{\bm{C}}$, such that $\tilde{\bm{A}}\bm{u}=\tilde{\bm{C}}\bm{f}$ has the same solution as $\bm{A}\bm{u}=\bm{C}\bm{f}$. We can derive them with the following transformations, by using the fact that $\bm{C}\bm{u}=\bm{C}\bm{f}$, as the inpainting solution and the original image coincide at the mask pixels:
@@
We note that one could also apply other solvers for symmetric problems to solve $\tilde{\bm{A}}\bm{u}=\tilde{\bm{C}}\bm{f}$, such as the conjugate residual (CR) method, the minimal residual (MINRES) solver [58], or the symmetric LQ (SYMMLQ) algorithm [58]. However, while those methods minimize the residual faster compared to conjugate gradients, and work for indefinite systems, CG remains the fastest w.r.t. minimizing the error to the solution $\bm{u}^{*}$ for our problem. In our case $\tilde{\bm{A}}$ is positive definite as long as $\bm{c}\neq\bm{0}$, so we leverage this fact by using CG.
@@
We improve the nested CG approach in two ways. First, we replace the inner CG solvers for the two inpainting problems by the multigrid ORAS inpainting solver, which improves the runtime by more than a factor of 4. Secondly, we propose to replace the outer CG solver with CGNR, a variant of CG for the normal equations [12, 37, 69]. Besides better numerical stability compared to CG, CGNR also has the advantage that the original inpainting residual $\bm{r}=\bm{f}-\bm{A}\bm{u}$, and hence the MSE, is available in each iteration. Because we want to optimize the MSE, a stopping criterion based on the relative improvement of the MSE in an iteration is better suited for our purposes than the standard stopping criterion based on the $2$-norm of the residual of the normal equations $\|\bm{B}^{\top}(\bm{f}-\bm{B}\bm{g})\|_{2}$.
@@ 5.2 RAS Solver for Tonal Optimization
Taking into account the above, we solve the normal equations with a domain decomposition method instead of CGNR. We adapt a restricted additive Schwarz (RAS) method [18], which is similar to the one used in the multigrid ORAS inpainting solver. Similar to the inpainting solver, we subdivide the image domain $\Omega$ into $k\in\mathbb{N}$ overlapping subdomains $\Omega_{1},...,\Omega_{k}\subset\Omega$, such that $\cup_{i=1}^{k}\Omega_{i}=\Omega$. The resulting RAS tonal optimization is given in Algorithm 2. It consists of three steps that are iterated until the method is sufficiently converged.
@@
    (18) $\bm{g}^{n+1}=\bm{g}^{n}+\sum_{i=1}^{k}\bm{R}_{i}^{\top}\bm{D}_{i}\bm{v}_{i}^{n}$
First we compute the global residual $\bm{r}^{n}$. This involves the application of the system matrix. As in Subsection 5.1 we evaluate the matrix-vector products without explicitly forming the matrix, by solving two inner inpainting problems. As the inpainting solution changes only slightly from one iteration to the next, it can be used as a good initialization for the next inpainting. This allows us to skip the initialization phase in the full multigrid ORAS inpainting solver and start directly with a V-cycle at the finest resolution level, which improves the runtime of both inpaintings significantly. This is an advantage of RAS compared to CGNR, where we apply the system matrix only to conjugate vectors, which change more from one iteration to the next, and thus cannot be used as a good initialization.
@@
Overall our RAS method consists of four types of problems, which are shown in Fig. 6. Our global tonal optimization problem is solved with RAS which includes two global inner inpainting problems per iteration to compute the global residual. As with our spatial optimization method, we solve them with the multigrid ORAS inpainting solver. For each domain decomposition block in the RAS solver we also get local tonal optimization problems which we solve with CGNR. These again include two local inner inpainting problems per iteration that we solve with the same multigrid ORAS inpainting method.
For the domain decomposition we use blocks of size $64\times 64$, an overlap of $6$ pixels and a simple averaging for mask pixels in the overlap. The size of the blocks should be a multiple of $32$ due to the architecture of our GPU in order to achieve a good performance. We found that a block size of $64$ and an overlap of $6$ pixels between them is optimal for our GPU, as the blocks are still quite small while the average number of mask pixels per block and the overlap is sufficiently large for a good convergence towards the global solution.
@@ 5.3 Voronoi Initialization
(19) $\bm{g}_{i}=\bm{u}_{i}+\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}(\bm{f}_{j}-\bm{u}_{j})\quad\text{for all }i\in K,$
@@
(20) $\bm{C}\bm{W}\bm{A}^{-1}\bm{C}\bm{g}=:\bm{C}\bm{W}\bm{u}=\bm{C}\bm{W}\bm{f}.$
We solve (20) with a modified Richardson iteration:
(21) $\bm{C}\bm{g}^{k+1}=\bm{C}\bm{g}^{k}+\tau(\bm{C}\bm{f}-\bm{C}\bm{W}\bm{A}^{-1}\bm{C}\bm{g}^{k}).$
@@ 6.1 Experimental Setup
For the experimental evaluation of our Delaunay densification spatial optimization method, we compare against the analytic approach (AA) by Belhachmi et al. [8], probabilistic sparsification (PS) [54], with (PS+NLPE) and without nonlocal pixel exchange, and the most recent neural approach by Schrader et al. [71]. The analytic approach is implemented with Floyd-Steinberg dithering [31] of the Laplacian magnitude. To ensure fair comparisons, PS and NLPE are both implemented on the GPU with the same multigrid ORAS inpainting solver that we use for our methods. For PS we use candidate fractions $p=0.3$ and $q=0.005$, and NLPE is run for five cycles. As the neural approach requires more than 20 GB of graphics memory for runtime tests, we exclude them in our runtime comparison.
@@ 6.2 Spatial Optimization Experiments
In Fig. 9(a) we evaluate the impact of the choice of the initial mask on our Delaunay densification. We compare a uniform random initial mask with a Laplacian magnitude sampled initial mask. We use the same number of initial mask pixels and add a constant amount of new pixels in each iteration for both masks. We observe that for a small amount of iterations the Laplacian magnitude approach results in a significantly higher PSNR, while the difference becomes smaller the more iterations are used. This can be explained with the fact that fewer mask pixels are introduced per iteration if we have a larger numbers of iterations, which reduces the impact of the initial mask. It also shows that the improvement per iteration becomes quite small for iterations beyond $20$, while the additional runtime for each iteration is approximately constant as each requires exactly one inpainting. Thus $20$ iterations seem to be a good trade-off between runtime and reconstruction quality.
@@
Figure 9(b) shows the effect of increasing or decreasing the number of mask pixels that are added per iteration during the densification with a constant factor $t$. For a uniform random initial mask we observe that adding a constant amount of pixels in each iteration is not the optimal strategy. By increasing the number of added pixels by a factor of $t=1.08$ or $8$% in each iteration, we can improve the PSNR by $0.05$ dB. With a sampled Laplacian magnitude as an initial mask, however, the optimal factor is only $t=1.03$ and the improvement compared to a constant number of pixels is less than $0.01$ dB. These results also reinforce the conclusion that the Laplacian magnitude yields a suitable initial mask and suggests only slightly increasing the number of pixels added per iteration. Since the difference between adding a constant number of pixels and the optimal strategy is negligible, for simplicity we use a constant number of pixels for the rest of our experiments.
(a) PSNR vs iterations
(b) PSNR vs factor
@@
We evaluate the runtimes in Fig. 10(b). While the inferior analytic approach is clearly the fastest, our method is the second fastest with an average runtime of around $0.4$ seconds. We achieve this by restricting our densification to $20$ iterations, which still produces masks of very high quality. Compared to PS and PS+NLPE we are several orders of magnitude faster, since they both require thousands of inpaintings. For the neural approach we cannot generate any runtime results on our GPU due to memory restrictions, but the runtimes reported by Schrader et al. [71] are quite similar to the results of our approach, albeit a bit faster for densities below $3$%. As they are obtained with a significantly more powerful GPU, our method should be faster on a similar GPU. Furthermore, our method has the advantage of needing significantly less GPU memory.
@@
The qualitative evaluation of the spatial optimization methods in Fig. 10(a) shows that with just $20$ densification iterations our Delaunay densification approach (DD) already consistently outperforms all its competitors on all tested mask densities. It even surpasses the recent neural approach by Schrader et al. [71] by at least $0.75$ dB. Compared to the model-based competitors our method is especially well-suited for lower mask densities, which are the most interesting for inpainting-based compression.
Here our method outperforms the analytic approach by more than $6$ dB, PS by more than $3.8$, and PS+NLPE by more than $2.6$ dB. This suggests that densification together with an adaptation to the local mask density with for example the Delaunay triangulation should be preferred over the probabilistic sparsification approach, as it significantly outperforms it in terms of quality as well as runtime.
@@ 6.3 Tonal Optimization Experiments
In order to quantify the effect of our Voronoi approach on the tonal optimization problem, we perform an experiment using our RAS tonal optimization method with and without the Voronoi initialization. Additionally, we compare against a CGNR solver with the same multigrid ORAS inpainting. Fig. 14(b) shows the result for a $5$% mask density. For a better visualization, we have omitted the results for the CGNR solver with CG inpainting, as each iteration needs approximately $0.8$ seconds, which makes it clearly inferior to the other methods. We can observe that RAS converges significantly faster to the optimal PSNR compared to the CGNR method. When we use RAS together with the Voronoi initialization, we can stop the tonal optimization much earlier, even though the initial iteration takes longer. As the tonal optimization is a convex least squares problem, all tested methods eventually converge to the same optimal PSNR. Therefore, we omit a comparison of the reconstruction quality, as all methods end up with the same solution. The experiment also shows that the Voronoi initialization alone almost reaches the optimal PSNR up to $0.2$ dB. This proves that the Voronoi initialization is an essential tool for tonal optimization.
@@ 7 Conclusions and Outlook
Inpainting-based compression does not only require a good inpainting method but also methods for the optimization of the inpainting data. The optimization of the spatial data as well as the tonal data for homogeneous diffusion inpainting is a challenging task, and previous approaches were either slow or of inferior quality. As a remedy, we developed new methods for the data optimization of homogeneous diffusion inpainting that outperform previous approaches by a wide margin. Furthermore, most prior approaches could handle only relatively small images, while we are able to work with images of 4K resolution and higher with optimal linear resolution to runtime scaling. We achieved this by carefully adapting some of the most successful numerical concepts and transferring them to an image compression application. For our spatial optimization method we combined a mask densification approach with error-map dithering ideas using a Delaunay triangulation. For the tonal optimization we adopted a domain decomposition method that solves the corresponding normal equations in a matrix-free fashion. Additionally we proposed a Voronoi-based initialization strategy that allows us to stop the tonal optimization much earlier. This allows us to create high quality inpainting masks and optimized tonal values in a runtime of less than a second on a contemporary GPU. This constitutes a significant improvement over any comparable previous approaches and suggests a similar order of runtime efficiency as JPEG2000, while solving a numerically much more challenging problem. We note, however, that the runtimes in our approach are for the data optimization and do not include an encoding step currently. Together with a real-time capable inpainting (see [49]), we presented a basis for inpainting-based compression methods that may serve as serious alternatives to classical transform-based codecs, both in terms of quality as well as runtimes.
So far we only considered the data optimization for homogeneous diffusion inpainting. However, many of our proposed methods can be extended to more sophisticated operators, such as anisotropic nonlinear diffusion that are able to qualitatively outperform homogeneous diffusion. Our Delaunay-based spatial optimization and our Voronoi-based initialization strategy for the tonal optimization are fairly general and can be used with most inpainting operators. In our ongoing research we investigate how well these methods perform with different inpainting operators and how they can be improved.